# Differential Privacy of Cross-Attention with Provable Guarantee

### v2

@@ Appendix E Weighted $\ell_{p}^{p}$ Distance
Therefore, we can solve one dimension problem first, and then the high dimension case can be solved automatically.
E.1 One Dimensional Weighted Distance
@@
Lemma E.1 (Weighted distance one dimension).
If the following conditions hold:
  • Let data $X\in[0,R]^{n}$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]$.
  • We round $X$ and $y$ to an integer multiple of $R/n$.
  • Let $c_{j}=\sum_{j_{0}\in S_{j}}w_{j_{0}}$ where set $S_{j}$ is the set of index $i$ such that the corresponding $x_{i}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$.
  • After rounding, we assume $y$ is in the $kR/n$ position for $k\in\{0,1,2,\dots,n\}$.
For the weighted problem, we have
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|=\sum_{j=0}^{n}(|k-j|R/n+O(R/n))c_{j}.$
Moreover, we have
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(|k-j|R/n+O(R/n))^{p}c_{j}$
where $O(R/n)$ is the rounding error for each data point.
@@
For each $i$, we have:
$w_{i}\cdot|y-x_{i}|=w_{i}\cdot(\frac{|k-j|R}{n}+O(\frac{R}{n})).$
where $O(R/n)$ is the rounding error introduced by each data point, since each data point will be at most $O(R/n)$ away from its true position.
We can construct $c_{j}$ by
$c_{j}=\sum_{j_{0}\in S_{j}}w_{j_{0}}$
set $S_{j}$ is the set of index $i$ such that the corresponding $x_{i}$ is rounded to $jR/n$. Moreover, $c_{j}$ can be negative.
Summing over all $i$ and grouping by $j$, we get:
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|=\sum_{j=0}^{n}(\frac{|k-j|R}{n}+O(\frac{R}{n}))c_{j}.$
The total rounding error will be $O(R)$ because we have $n$ data points, each with an error of at most $O(R/n)$.
Moreover, we have
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(\frac{|k-j|R}{n}+O(\frac{R}{n}))^{p}c_{j}.$
E.2 High Dimensional Weighted Distance
Finally, we can solve the problem of weighted distance for $d$-dimensional dataset.
Lemma E.2 (Weighted $\ell_{p}^{p}$-distance high dimension, formal version of Lemma C.4).
If the following conditions hold:
  • Let data $X\in[0,R]^{n\times d}$ and $x_{i}^{\top}\in[0,R]^{d}$ be the $i$-th row of $x$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]^{d}$.
  • We round each dimension of $X$ and $y$ to an integer multiple of $R/n$.
  • Let $x_{i,k},y_{k}$ denote the $k$-th coordinates of $x_{i},y$ for $k\in[d]$.
  • Let $c_{j,k}:=\sum_{j_{0}\in S_{j,k}}w_{j_{0}}$ where set $S_{j,k}$ is the set of index $i$ such that the corresponding $x_{i,k}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
  • After rounding, we assume $y_{k}$ is in the $l_{k}R/n$ position for $l_{k}\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
For the weighted problem, we have
$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}$
where $O(R/n)$ is the rounding error for each data point.
Proof.
We can show
$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=$ $~{}\sum_{k=1}^{d}\sum_{i=1}^{n}w_{i}\cdot|y_{k}-x_{i,k}|^{p}$
$=$ $~{}\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}$
where the first step follows from decomposability of $\ell_{p}^{p}$-distance by dimension, the second step follows from Lemma E.1.
Appendix F One-Dimensional Weighted $\ell_{1}$ Distance Query
In this section, we generalize the algorithms in BLM+ [10] to weighted distance. Here, we compute the problem of one-dimensional weighted $\ell_{1}$ distance query i.e. $\sum_{i\in[n]}w_{i}|y-x_{i}|$ for a given query $y\in[0,R]$, weights $w\in[-R_{w},R_{w}]^{n}$ and dataset $X\subset[0,R]$ and $n=|X|$. In Section F.1, we analyze the runtime of our algorithm. In Section F.2, we analyze the DP and accuracy of our algorithm. In Section F.3, we give the theorem for our DPTreeDistance data structure.
@@ Appendix F One-Dimensional Weighted $\ell_{1}$ Distance Query
1:datastructure DPTreeDistance $\triangleright$ Theorem F.6
@@
3:     ${\mathcal{D}}:$ DPTree $\triangleright$ Alg. 2
@@
7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma E.1
8:     $X,w,a\leftarrow X,w,0^{n+1}$
9:     for $i=1\to n$ do
10:         $j\leftarrow\textsc{Round}(x_{i},n)$triangleright $x_{i}\in X$ for $i\in[n]$
11:         $a_{j}\leftarrow a_{j}+w_{i}$
12:     end for
13:     ${\mathcal{D}}$.$\textsc{Init}(a,n+1,2R_{w},\epsilon,\delta)$triangleright Alg. 2, Lemma C.3
14:end procedure
15:procedure Round($x\in[0,R]$, $n\in\mathbb{N}_{+}$)
16:     Let $j\in\{0,1,2,\dots n-1\}$ denote the integer such that $jR/n\leq x<(j+1)R/n$
17:     if $|x-(j+1)R/n|\leq|x-jR/n|$ then
18:         $j\leftarrow j+1$
19:     end if
20:     return $j$
21:end procedure
22:end datastructure
Algorithm 6 One dimensional weighted $\ell_{1}$ distance query
1:datastructure DPTreeDistance $\triangleright$ Theorem F.6
2:procedure DistanceQuery($y\in[0,R]$, $\alpha\in(0,1)$) $\triangleright$ Lemma F.2, Lemma F.4, Lemma F.5
3:     $y\leftarrow\textsc{Round}(y,n)\cdot(R/n)$triangleright Alg. 5
4:     Value $\leftarrow 0$
5:     for $j=0,1,...,O(\log(n)/\alpha)$ do
6:         $l_{j}\leftarrow\textsc{Round}(y+\frac{R}{(1+\alpha)^{j+1}},n)$
7:         $r_{j}\leftarrow\textsc{Round}(y+\frac{R}{(1+\alpha)^{j}},n)$triangleright Consider the points to the right of $y$
8:         Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$triangleright Alg. 2
9:     end for
10:     for $j=0,1,...,O(\log(n)/\alpha)$ do
11:         $l_{j}\leftarrow\textsc{Round}(y-\frac{R}{(1+\alpha)^{j}},n)$
12:         $r_{j}\leftarrow\textsc{Round}(y-\frac{R}{(1+\alpha)^{j+1}},n)$triangleright Consider the points to the left of $y$
13:         Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$triangleright Alg. 2
@@
15:     Return Value
16:end procedure
17:end datastructure
F.1 Runtime Analysis
We first analyze the runtime.
Lemma F.1 (Runtime of initialization, Algorithm 5).
For the initialization, we have the time complexity of Init in Algorithm 5 is $O(n)$.
Proof.
In the initialization of Init, the computations need $O(n)$ time to compute the count and $O(\log n)$ time to build the tree. Thus, total time is $O(n)$. ∎
Lemma F.2 (Runtime of DistanceQuery, Algorithm 6).
For the $\ell_{1}$ distance query, we have the time complexity of DistanceQuery in Algorithm 6 is $O(\alpha^{-1}\log^{2}n)$.
Proof.
In DistanceQuery, the computations need $O(\log n)$ time to compute one value from DPTree.Query and this process need to be repeated $O(\alpha^{-1}\log n)$ times. ∎
Remark F.3.
In Line 8 and 13 of Algorithm 6, we use $R/(1+\alpha)^{j}$ to approximate the distance of each data point to the query in Lemma E.1, i.e. $|k-j|R/n$. This will introduce $\alpha$ relative error but also reduce the numbers of iteration from $O(n)$ to $O(\log(n)/\alpha)$.
F.2 Privacy and Accuracy Analysis
We show the DP.
Lemma F.4 (Privacy of DistanceQuery, Algorithm 6).
The output process of DistanceQuery (Algorithm 6) is $(\epsilon,\delta)$-DP.
Proof.
From Lemma D.5, we know that $\mathrm{DPTree}$.Query is $(\epsilon,\delta)$-DP output. We observe that intervals in Algorithm 6 are disjoint. Then, following the same logic in the proof of Lemma D.8, we can conclude that the value returned is $(\epsilon,\delta)$-DP. ∎
We now analyze the accuracy of the algorithm.
Lemma F.5 (Accuracy of DistanceQuery, Algorithm 6).
If the following conditions are satisfied:
  • Let $X\in[0,R]^{n}$ be a dataset consisting of $n$ one-dimensional numbers, with weights $w\in[-R_{w},R_{w}]^{n}$.
  • Let $\alpha\in(0,1)$ represent the relative error parameter utilized in Algorithm 6.
  • Let $\widetilde{A}$ denote the output of the DistanceQuery in Algorithm 6.
  • Let $A_{*}:=\sum_{i\in[n]}w_{i}|y-x_{i}|$ represent the true distance query value for a specific query $y$.
Then with probability $0.99$, we have
$|\widetilde{A}-A_{*}|\leq\alpha A_{*}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n).$
Proof.
@@ F.2 Privacy and Accuracy Analysis
$\widehat{A}:=$ $~{}\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot\textsc{TrueQuery}(S_{j}).$
Since TrueQuery returns the sum of the corresponding weights, it aligns with the true answer $A_{*}:=\sum_{i\in[n]}w_{i}|y-x_{i}|$. Thus, we have
$|\widehat{A}-A_{*}|\leq\alpha\cdot A_{*},$
because for all $j$, the distances between $y$ and different points in $X$ vary only by a multiplicative factor of $(1+\alpha)$.
Next we show the additive error. Let $\textsc{Query}(S_{j})$ denote the noised interval query answer returned by DPTree.Query in Algorithm 2. Algorithm 6 outputs $\widetilde{A}=\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot\textsc{Query}(S_{j})$. We wish to bound
$|\widehat{A}-\widetilde{A}|\leq$ $~{}|\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot(\textsc{TrueQuery}(S_{j})-\textsc{Query}(S_{j}))|.$
Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$, which from Algorithm 2 we can see this is the sum of $O(\log n)$ independent truncated Laplace random variables.
From Lemma D.8, we only need to show that the series $\frac{1}{(1+\alpha)^{j}}$ for $j\in\{0,1,\dots,O(\log(n)/\alpha)\}$ square converges to $1/\alpha$, since $R$ is a constant.
We can show
$\sum_{j=0}^{O(\log(n)/\alpha)}\frac{1}{(1+\alpha)^{2j}}\leq$ $~{}\sum_{j=0}^{\infty}\frac{1}{(1+\alpha)^{2j}}$
$\leq$ $~{}\sum_{j=0}^{\infty}\frac{1}{(1+\alpha)^{j}}$
$=$ $~{}\frac{1}{1-\frac{1}{1+\alpha}}$
$=$ $~{}1+\frac{1}{\alpha}$
$=$ $~{}O(1/\alpha)$
where the first step follows from we extend the finite sum to infinite sum, the second step follows from $\frac{1}{(1+\alpha)^{2j}}\leq\frac{1}{(1+\alpha)^{j}}$, the third step follows from the closed form of geometric sum, the fourth step follows from simple algebra, and the last step follows from $\alpha\in(0,1)$.
Then from the proof of Lemma D.8, we can know that the variance is given by
$O(\frac{R^{2}R_{w}^{2}\log^{3}n}{\alpha\epsilon^{2}})$ (5)
since the sensitivity $\Delta=2R_{w}$ from Lemma C.3.
Using Lemma B.3, we can have additive error bounded by
$O(\frac{R\cdot R_{w}\log^{3/2}n}{\epsilon\sqrt{\alpha}}).$
with probability $0.99$. ∎
F.3 One Dimension Single Data Structure
We therefore have the data structure that can solve weighted $\ell_{1}$-distance problem.
Theorem F.6 (DPTreeDistance data structure, formal version of Theorem C.6).
There is a data structure DPTreeDistance (Algorithm 5,6) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 6) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
@@ F.3 One Dimension Single Data Structure
The proofs follow from combining Lemma F.1 (running time of initialization), Lemma F.2 (running time of query), Lemma F.4 (DP of query), and Lemma F.5 (error of query) together. ∎
Appendix G High-Dimensional Weighted $\ell_{1}$ Query
In this section, we show how we can solve the high dimensional weighted $\ell_{1}$ distance problem, generalizing results from BLM+ [10]. In Section G.1, we give the analysis of Algorithm 7. In Section G.2, we give the theorem of our DPTreeHighDim data structure.
Algorithm 5,6 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{1}$ distance function. We construct $d$ separate one-dimensional distance query data structures, each corresponding to a coordinate projection of the dataset.
G.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
We now give the analysis of our Algorithm 7 for high dimensional weighted $\ell_{1}$-distance query.
Algorithm 7 High-dimensional weighted $\ell_{1}$ distance query
1:datastrucutre DPTreeHighDim $\triangleright$ Theorem G.3

### v3

@@ Appendix C Weighted $\ell_{p}^{p}$ Distance
Lemma C.1 (Weighted distance one dimension).
For a collection of numbers $\{x_{1},x_{2},\cdots,x_{n}\}\subset\mathbb{R}$ and corresponding weights $\{w_{1},w_{2},\cdots,w_{n}\}\subset\mathbb{R}$, and a number $y\in\mathbb{R}$. We define two sets
$S_{+}:=$ $~\{k\in[n]~:~x_{k}>y\}$
$S_{-}:=$ $~\{k\in[n]~:~x_{k}<y\},$
It holds
$\sum_{k=1}^{n}w_{k}|x_{k}-y|^{p}=\sum_{j=0}^{p}{p\choose j}y^{p-j}((-1)^{p-j}\sum_{k\in S_{+}}w_{k}x_{k}^{j}+(-1)^{j}\sum_{k\in S_{-}}w_{k}x_{k}^{j}),$
where $p\choose j$} denotes the binomial coefficient that ${p\choose j}=\frac{p!}{j!(p-j)!}$.
@@
We show that
$\sum_{k=1}^{n}w_{k}|x_{k}-y|^{p}=$ $~\sum_{x_{k}\in S_{+}}w_{k}(x_{k}-y)^{p}+\sum_{x_{k}\in S_{-}}w_{k}(y-x_{k})^{p}$
$=$ $~(\sum_{x_{k}\in S_{+}}w_{k}\sum_{j=0}^{p}(-1)^{p-j}{p\choose j}x_{k}^{j}y^{p-j})+(\sum_{x_{k}\in S_{-}}w_{k}\sum_{j=0}^{p}(-1)^{j}{p\choose j}x_{k}^{j}y^{p-j})$
$=$ $~\sum_{j=0}^{p}({p\choose j}(-1)^{p-j}y^{p-j}\sum_{k\in S_{+}}w_{k}x_{k}^{j})+\sum_{j=0}^{p}({p\choose j}(-1)^{j}y^{p-j}\sum_{k\in S_{-}}w_{k}x_{k}^{j})$
$=$ $~\sum_{j=0}^{p}{p\choose j}y^{p-j}((-1)^{p-j}\sum_{k\in S_{+}}w_{k}x_{k}^{j}+(-1)^{j}\sum_{k\in S_{-}}w_{k}x_{k}^{j}).$
Thus, we complete the proof. ∎
Appendix D One-Dimensional Weighted $\ell_{p}^{p}$ Distance Query
In this section, we generalize the algorithms in [8] and [45] to weighted distance. Here, we compute the problem of one-dimensional weighted $\ell_{p}^{p}$ distance query i.e. $\sum_{i\in[n]}w_{i}|y-x_{i}|$ for a given query $y\in[0,R]$, weights $w\in[-R_{w},R_{w}]^{n}$ and dataset $X\subset[0,R]$ and $n=|X|$. In this section, we give the theorem for our DPTreeDistance data structure.
@@ Appendix D One-Dimensional Weighted $\ell_{p}^{p}$ Distance Query
1:datastructure DPTreeDistance $\triangleright$ Theorem D.1
@@
3:   ${\mathcal{D}}_{0},\dots,{\mathcal{D}}_{p}:$ DPTree $\triangleright$ Alg. 2
@@
7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma C.1
8:  $X,w,a\leftarrow X,w,0^{n\times(p+1)}$
9:  for $i=1\to n$ do $\triangleright$ $x_{i}\in X$ for $i\in[n]$
10:   Let $j\in[n]$ denotes the integer such that $x_{i}\in[(j-1)R/n,jR/n)$
11:   for $q=0\to p$ do
12:     $a_{j,q}\leftarrow a_{j,q}+w_{i}x_{i}^{q}$
13:   end for
@@
15:  for $q=0\to p$ do
16:   ${\mathcal{D}}_{q}$$\textsc{Init}(a_{:,q},n,2R_{w}R^{q},\epsilon/(p+1),\delta/(p+1))$ $\triangleright$ Alg. 2, Lemma B.3
17:  end for
18:end procedure
19:end datastructure
Algorithm 5 One dimensional weighted $\ell_{p}^{p}$
distance query
1:datastructure DPTreeDistance $\triangleright$ Theorem D.1
2:procedure DistanceQuery($y\in[0,R]$)
3:  for $q=0\to p$ do
4:   $c_{\mathrm{left},q},c_{\mathrm{right},q}\leftarrow$ $\mathcal{D}_{q}.$Query($y$)
5:  end for
6:  return $\sum_{q=0}^{p}{p\choose q}y^{p-q}((-1)^{p-q}c_{\mathrm{right},q}+(-1)^{q}c_{\mathrm{left},q})$
7:end procedure
8:end datastructure
Theorem D.1 (DPTreeDistance data structure ).
There is a data structure DPTreeDistance (Algorithm 4,5) that uses $O(np)$ spaces to solve weighted $\ell_{p}^{p}$ distance query problem for dataset $X\subset[0,R]$ and support the following operations:
@@
[…]
@@
We set the total layers of one tree $L=(\log n)$. There are $p+1$ trees.
Init Time and Space.
The total number of nodes on one tree is $O(n)$. There are total $O(pn)$ values stored for $p+1$ trees. Adding the time of iterating all data points, initializing these values takes $O(pn)$ time.
Query Time.
Each query iterates through all layers. On each layer it takes $O(1)$ time to calculate $c_{\text{left},q}$ and $c_{\text{right},q}$. There are $(\log n)$ layers, and $p+1$ trees, so the total query time is $O(p\log n)$.
Privacy Guarantees.
For each $\mathcal{D}_{q}$ for $q\in\{0,1,\dots,p\}$, we input $a_{:,q}$. Since $X\in[0,R]^{n}$ and $w\in[-R_{w},R_{w}]^{n}$, the input range for $a_{:,q}$ is $R_{w}R^{q},R_{w}R^{q}]$. Then from Lemma B.3, sensitivity is $2R_{w}R^{q}$.
From Lemma B.6, we know each $\mathcal{D}_{q}$ query is $(\epsilon/(p+1),\delta/(p+1))$-DP. By basic composition Fact A.8, the total differential privacy parameter is $(\epsilon,\delta)$. This completes the proof.
Error Guarantees.
The additive error consists of two parts.
The first part is from the data in the leaf node which contains query $y$. The error is
$\sum_{x_{k}\in[(j-1)\cdot R/2^{L},j\cdot R/2^{L})}|x_{k}-y|^{p}\leq n\cdot(\frac{R}{2^{L}})^{p}.$
When $L=\log n$, this error is $O(R^{p}/n^{p-1})$.
The second part is the Truncated Laplace noise. From the proof of Lemma B.7, we have each $\mathcal{D}_{q}$ for $q\in\{0,1,\dots,p\}$ has $O(L)$ independent $\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)$ noises for $L=\log n$ layers.
Let $A$ be the noisy output of DistanceQuery in Algorithm 5 and $A_{*}=\sum_{k\in[n]}w_{k}|y-x_{k}|$ be the true output. Then, for our Algorithm 4 and 5, the variance is
$\operatorname*{\mathrm{Var}}[\sum_{i}^{L}\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)]=$ $\sum_{i}^{L}\operatorname*{\mathrm{Var}}[\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)]$
$=$ $~O(L^{3}\epsilon_{q}^{-2}\Delta_{q}^{2})$
Replacing $\Delta_{q}=O(R^{q}R_{w})$ and $\epsilon_{q}=O(\epsilon/p)$, using Lemma A.3, with high probability 0.99, we have
$|\sum_{i}^{L}\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)|\leq O(pR_{w}R^{q}L^{3/2}/\epsilon).$ (5)
Then we bound the error with this inequality:
$|A-A^{\prime}|\leq$ $~|\sum_{q=0}^{p}{p\choose q}y^{p-q}\sum_{i=1}^{L}((-1)^{p-q}\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)+(-1)^{q}\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L))|$
$\leq$ $~\sum_{q=0}^{p}{p\choose q}y^{p-q}|\sum_{i=1}^{L}(\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)+\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L))|$
$=$ $~\sum_{q=0}^{p}{p\choose q}y^{p-q}\cdot O(pR_{w}R^{q}L^{3/2}/\epsilon)$
$=$ $~O(\epsilon^{-1}pR_{w}L^{3/2}\sum_{q=0}^{p}{p\choose q}y^{p-q}R^{q})$
$=$ $~O(\epsilon^{-1}pR_{w}L^{3/2}(y+R)^{p})$
$=$ $~O(\epsilon^{-1}pR_{w}(2R)^{p}\log^{3/2}n),$
where the third step follows from Eq. (5), and the last step is from $L=\log n$ and $y\in[0,R]$.
Therefore, by triangle inequality and two parts of error, the total error is
$O(R^{p}/n^{p-1})+O(\epsilon^{-1}pR_{w}(2R)^{p}\log^{3/2}n)\leq O(\epsilon^{-1}pR_{w}(2R)^{p}\log^{3/2}n),$
since $p\geq 1$ and $n\in\mathbb{N}_{+}$. This completes the proof. ∎
Appendix E High-Dimensional Weighted $\ell_{p}^{p}$ Query
In this section, we show how we can solve the high dimensional weighted $\ell_{p}^{p}$ distance problem, generalizing results from [8] and [45]. In Section E.1, we give the analysis of Algorithm 6. In Section E.2, we give the theorem of our DPTreeHighDim data structure.
Algorithm 4,5 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{p}^{p}$ distance function. We construct $d$ separate one-dimensional distance query data structures, each corresponding to a coordinate projection of the dataset.
E.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
We now give the analysis of our Algorithm 6 for high dimensional weighted $\ell_{p}^{p}$ distance query.
Algorithm 6 High-dimensional weighted $\ell_{p}^{p}$ distance query
1:datastrucutre DPTreeHighDim $\triangleright$ Theorem E.3